What is the capacitance of a capacitor that stores 50 picocoulomb of charge when connected to a 4 V battery?


1
Expert's answer
2022-03-11T08:28:33-0500

The capacitance is given as follows:


"C = \dfrac{Q}{\Delta V}"

where "Q = 50\times 10^{-12}C, \Delta V = 4V". Thus, obtain:


"C = \dfrac{50\times 10^{-1}C}{4V} = 12.5\times 10^{-12}F"

Answer. "12.5\times 10^{-12}F".